OpenStudy (anonymous):

help?!? A circle has its center at (-2, 5) and a radius of 4 units. What is the equation of the circle? (x + 2)2 + (y - 5)2 = 16 (x + 2)2 + (y + 5)2 = 16 (x + 2)2 + (y - 5)2 = 4 (x - 2)2 + (y + 5)2 = 4

jigglypuff314 (jigglypuff314):

the equation of a circle: (x - h)^2 + (y - k)^2 = r^2 (h, k) = center and r = radius so for when its center is at (-2, 5) and it has a radius of 4 plug in h=-2 , k=5 , and r=4 into the circle equation

OpenStudy (anonymous):

the answer will be D

jigglypuff314 (jigglypuff314):

I don't think the answer would be D... because of the r^2 part 4^2 = 16

OpenStudy (anonymous):

No. not d. Look at Jiggly's formla.....h is -2....so you get (x - (-2))^2 in the formula. also k is 5 ... so you get (y - 5)^2 in the formula

OpenStudy (anonymous):

will it be A

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

thx

jigglypuff314 (jigglypuff314):

you're welcome :)